One canned juice drink 30% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 25L that is 20% orange juice.
How many liters of the 30% orange juice should be in the mixture? _L
How many liters of the 5% orange juice should be in the mixture?_L

Answer by Alan3354(69443)

One canned juice drink 30% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 25L that is 20% orange juice.
-------------
t = amount of 30%
f = amount of 5%
----
t + f = 25
30t + 5f = 25*20